# Visual Generation Without Guidance

### v1

@@ 3.3 GFT for AR and Masked Models
CFG is also a standard decoding method for discrete AR or masked visual models, critical to improving their sample quality (Chang et al., 2022; Li et al., 2023; Tian et al., 2024). Different from diffusion models which apply CFG in the score field, AR and masked models perform guided sampling on model logits:
$\ell_{\theta}^{\text{s}}({\bm{x}}_{n}|{\bm{x}}_{<n},{\bm{c}})=$
Here ${\bm{x}}_{n}$ represents the $i$-th token of a tokenized image ${\bm{x}}$. $\ell_{\theta}$ are unnormalized model logits.
@@
${\mathcal{L}}_{\theta}^{\text{AR}}({\bm{x}},{\bm{c}}_{\varnothing},\beta)$ $=-\sum_{i}\log p_{\theta}^{s}({\bm{x}}_{n}|{\bm{x}}_{<n},{\bm{c}}_{\varnothing},\beta)$ (10)
$=-\sum_{i}\log\frac{e^{\ell_{\theta}^{s}({\bm{x}}_{n}|{\bm{x}}_{<n},{\bm{c}}_{\varnothing},\beta)}}{\sum_{w\in\mathcal{V}}e^{\ell_{\theta}^{s}(w|{\bm{x}}_{<n},{\bm{c}}_{\varnothing},\beta)}},$ (11)
@@
$\ell_{\theta}^{s}(w|{\bm{x}}_{<n},{\bm{c}}_{\varnothing},\beta)$
@@ 4 Connection with Other Guidance-Free Methods
CCA proves the optimal solution for solving Eq. 14 is $r_{\theta}^{*}=\log\frac{p({\bm{x}}|{\bm{c}})}{p({\bm{x}})}$, thus the convergence point for $p_{\theta}^{s}({\bm{x}}|{\bm{c}})$ also satisfies Eq. 5.
@@
Table 3: Zero-shot text-to-image generation performance comparison on COCO 2014 validation set.
@@
Stable Diffusion 1.5  (Rombach et al., 2022) 22.55 8.30
$\quad$++GFT (Ours) 8.32
@@ 5 Experiments
    GFT’s capacity in controlling diversity-fidelity trade-off through temperature parameter $\beta$. (Sec. 5.4)
w/o Guidance
@@ 5.1 Experimental Setups
We train C2I models on ImageNet-256x256 (Deng et al., 2009). For T2I models, we use a subset of the LAION-Aesthetic $5+$ (Schuhmann et al., 2022), consisting of 12.8 million image-text pairs. Our codebases are directly modified from the official CFG implementation of each respective baseline, keeping most hyperparameters consistent with CFG training. We use official OPENAI evaluation scripts to evaluate our C2I models. For T2I models, we evaluate our model on zero-shot COCO 2014 (Lin et al., 2014). The training and evaluation details for each model can be found in Appendix D.
@@ 5.2 Make CFG Models Guidance-Free
GFT achieves comparable performance to guidance distillation (designed for diffusion models) while outperforming AR alignment method in both performance (Table 2) and efficiency (Table 1). We attribute this effectiveness to GFT’s end-to-end training style, and to the nice convergence property of its maximum likelihood training objective.
@@ 5.3 Building Guidance-Free Models from Scratch
Figure 6: Comparison of training dynamics between GFT and CFG on diffusion and autoregressive models.
@@
Figure 8: FID-CLIP trade-off comparison on COCO-2014 validation set. Methods compared using Stable Diffusion 1.5 (SD 1.5).
@@ 5.4 Sampling Temperature for Visual Generation
We evaluate diversity-fidelity trade-offs across various models, with FID-IS trade-off for c2i models and FID-CLIP trade-off for t2i models. Results for DiT-XL/2 (fine-tuning), LlamaGen-L (from-scratch training) and Stable Diffusion 1.5 (fine-tuning) are shown in Figures 7 and 8, with additional trade-off curves provided in Appendix C. For DiT-XL/2, we also compare GFT with Guidance Distillation, showing that GFT achieves results comparable to both CFG and Guidance Distillation in the diversity-fidelity trade-off.
@@ 6 Conclusion
Acknowledgements
We thank Huanran Chen, Xiaoshi Wu, Cheng Lu, Fan Bao, Chengdong Xiang, Zhengyi Wang, Chang Li and Peize Sun for the discussion.
@@ Appendix A Related Work
Visual generative modeling has witnessed significant advancements in recent years. Recent explicit-likelihood approaches can be broadly categorized into diffusion-based models (Ho et al., 2020; Song et al., 2020b; Dhariwal & Nichol, 2021; Kingma et al., 2021; Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Karras et al., 2022; Bao et al., 2023; Peebles & Xie, 2023; Esser et al., 2024; Xie et al., 2024), auto-regressive models (Chen et al., 2020; Esser et al., 2021; Ramesh et al., 2021; Yu et al., 2021; Tian et al., 2024; Team, 2024; Sun et al., 2024; Ma et al., 2024; Zhang et al., 2024; Tang et al., 2024), and masked-prediction models (Chang et al., 2022; Yu et al., 2023a; Chang et al., 2023; Li et al., 2024; Yu et al., 2024; Fan et al., 2024). The introduction of guidance techniques has substantially improved the capabilities of these models. These include classifier guidance (Dhariwal & Nichol, 2021), classifier-free guidance (Ho & Salimans, 2022), energy guidance (Chung et al., 2022; Zhao et al., 2022; Lu et al., 2023; Song et al., 2023), and various advanced guidance methods (Kynkäänniemi et al., 2024; Karras et al., 2024; Chung et al., 2024; Koulischer et al., ; Shenoy et al., 2024).
@@
To address the computational overhead introduced by classifier-free guidance (CFG), One widely used approach to remove CFG is guidance distillation (Meng et al., 2023), where a student model is trained to directly learn the output of a pre-trained teacher model that incorporates guidance. This idea of guidance distillation has been widely adopted in methods aimed at accelerating diffusion models (Luo et al., 2023; Yin et al., 2024b, a; Zhou et al., 2024). By integrating the teacher model’s guided outputs into the training process, these approaches achieve efficient few-step generation without guidance.
Condition Contrastive Alignment.
Beyond guidance distillation, Condition Contrastive Alignment (CCA) (Chen et al., 2024b) could also significantly enhance the guidance-free performance of visual AR models through alignment (Rafailov et al., 2023; Chen et al., 2024a) in a self-contrastive manner.
@@ Appendix B Proof of Theorem 1
We first copy the training objective in Eq. 3.2 and Eq. 8.
@@
Given unlimited model capacity and training data, the optimal $\epsilon_{\theta^{*}}^{s}$ for optimizing Eq. 15 and Eq. 16 are the same. Both satisfy
@@
which is standard unconditional diffusion loss. According to Eq. 2 we have
@@
Then we prove stopping the unconditional gradient does not change this optimal solution. Taking derivatives of ${\mathcal{L}}_{\theta}^{\text{practical}}$ we have:
@@
Since $[2\mathbb{E}\beta]$ is a constant, this does not change the convergence point of ${\mathcal{L}}_{\theta}^{\text{raw}}$. The optimal unconditional solution for ${\mathcal{L}}_{\theta}^{\text{practical}}$ remains the same.
@@
We mainly apply GFT to fine-tune DiT-XL/2 (28 epochs, 2% of pretraining epochs) and train DiT-B/2 from scratch (80 epochs, following the original DiT paper’s settings (Peebles & Xie, 2023)). Since the DiT-B/2 pretraining checkpoint is not publicly available, we reproduce its pretraining experiment. For all experiments, we use a batch size of 256 and a learning rate of $41e-$ 4. For DiT-XL/2 fine-tuning experiments, we employ a cosine-decay learning rate scheduler.
@@
We mainly apply GFT to fine-tune VAR-d30 models (15 epochs) or train VAR-d16 models from scratch (200 epochs). Batch size is 768. The initial learning rate is $51e-$ 5 in fine-tuning experiments and $41e-$ 4 in pretraining experiments. Following VAR (Tian et al., 2024), we employ a learning rate scheduler including a warmup and a linear decay process (minimal is 1% of the initial).
VAR by default adopts a pyramid CFG technique on predicted logits. The guidance scale 00 decreases linearly during the decoding process. Specifically, let n be the current decoding step index, and N be the total steps. The $n$-step guidance scale $s_{n}$ is
@@
We mainly apply GFT to fine-tune LlamaGen-3B models (15 epochs) or train LlamaGen-L models from scratch (300 epochs). For fine-tuning, the batch size is 256, and the learning rate is $42e-$ 4. For pretraining, the batch size is 768, and the learning rate is $41e-$ 4. We adopt a cosine-decay learning rate scheduler in all experiments.
@@
We apply GFT to MAR-B, including both fine-tuning (10 epochs) and training from scratch (800 epochs). We find the batch size crucial for MAR and use 2048 following the original paper. For fine-tuning, we employ a learning rate scheduler including a 5-epoch linear warmup to $48e-$ 4 and a cosine decay process to $41e-$ 4. For training from scratch, we employ a 100-epoch linear lr warmup to $48e-$ 4, followed by a constant lr schedule, which is the same configuration as the original MAR pretraining.
@@
We apply GFT to fine-tune Stable Diffusion 1.5 (SD1.5) (Rombach et al., 2022) for 50,000 gradient updates with a batch size of 256 and constant learning rate of $51e-$ 5 with 1,000 warmup steps. We disable conditioning dropout as we find it improves CLIP score.
For evaluation, following GigaGAN (Kang et al., 2023) and DMD (Yin et al., 2024b), we generate images using 30K prompts from the COCO2014 (Lin et al., 2014) validation set, downsample them to 256×256, and compare with 40,504 real images from the same validation set. We use clean-FID (Parmar et al., 2022) to calculate FID and OpenCLIP-G (Ilharco et al., 2021; Cherti et al., 2023) to calculate CLIP score (Radford et al., 2021). All results are evaluated using 50 steps DPM-solver++ (Lu et al., 2022) with EMA models. The EMA decay rate is set to 0.999 for insufficient training steps.

### v2

@@ 3.3 GFT for AR and Masked Models
${\mathcal{L}}_{\theta}^{\text{AR}}({\bm{x}},{\bm{c}}_{\varnothing},\beta)$ $=-\sum_{i}\log p_{\theta}^{c}({\bm{x}}_{n}|{\bm{x}}_{<n},{\bm{c}}_{\varnothing},\beta)$ (10)
$=-\sum_{i}\log\frac{e^{\ell_{\theta}^{c}({\bm{x}}_{n}|{\bm{x}}_{<n},{\bm{c}}_{\varnothing},\beta)}}{\sum_{w\in\mathcal{V}}e^{\ell_{\theta}^{c}(w|{\bm{x}}_{<n},{\bm{c}}_{\varnothing},\beta)}},$ (11)
@@
$\ell_{\theta}^{c}(w|{\bm{x}}_{<n},{\bm{c}}_{\varnothing},\beta)$
@@ 4 Connection with Other Guidance-Free Methods
CCA proves the optimal solution for solving Eq. (14) is $r_{\theta}^{*}=\log\frac{p({\bm{x}}|{\bm{c}})}{p({\bm{x}})}$, thus the convergence point for $p_{\theta}^{s}({\bm{x}}|{\bm{c}})$ also satisfies Eq. (5).
@@
Table 3: Model comparisons for zero-shot text-to-image generation on the COCO 2014 validation set.
@@
Stable Diffusion 1.5  (Rombach et al., 2022) 22.55 7.87
[…]
@@ 5 Experiments
    GFT’s capability of controlling diversity-fidelity trade-off through temperature parameter $\beta$. (Sec. 5.4)
Vanilla conditional generation (w/o CFG)
w/ CFG
@@ 5.1 Experimental Setups
We train C2I models on ImageNet-256x256 (Deng et al., 2009). For T2I models, we use a subset of the LAION-Aesthetic $5+$ (Schuhmann et al., 2022), consisting of 18 million image-text pairs. Our codebases are directly modified from the official CFG implementation of each respective baseline, keeping most hyperparameters consistent with CFG training. We use official OPENAI evaluation scripts to evaluate our C2I models. For T2I models, we evaluate our model on zero-shot COCO 2014 (Lin et al., 2014). The training and evaluation details for each model can be found in Appendix D.
@@ 5.2 Make CFG Models Guidance-Free
GFT achieves comparable performance to guidance distillation (designed for diffusion models) and outperforms AR alignment method (Table 2 and 3). Notably, GFT demonstrates superior efficiency compared to both methods (Table 1). We attribute this effectiveness to GFT’s end-to-end training style, and to the nice convergence property of its maximum likelihood training objective.
@@ 5.3 Building Guidance-Free Models from Scratch
Figure 6: Comparison of convergence speed between GFT and CFG on diffusion and autoregressive models in from-scratch training.
@@
Figure 8: FID-CLIP trade-off comparison on COCO-2014 validation set. Methods compared using Stable Diffusion 1.5.
@@ 5.4 Sampling Temperature for Visual Generation
We evaluate diversity-fidelity trade-offs across various models, with FID-IS trade-off for c2i models and FID-CLIP trade-off for t2i models. Results for DiT-XL/2 (fine-tuning), LlamaGen-L (from-scratch training) and Stable Diffusion 1.5 (fine-tuning) are shown in Figures 7 and 8, with additional trade-off curves provided in Appendix C. For DiT-XL/2 and Stable Diffusion 1.5, we also compare GFT with Guidance Distillation, showing that GFT achieves results comparable to CFG while outperforming Guidance Distillation on CLIP scores.
@@ 6 Conclusion
Acknowledgement
We thank Huanran Chen, Xiaoshi Wu, Cheng Lu, Fan Bao, Chengdong Xiang, Zhengyi Wang, Chang Li and Peize Sun for the discussion. This work was supported by the NSFC Project (No. 62376131, 92270001, 92370124, 92248303) and the High Performance Computing Center, Tsinghua University. J.Z is also supported by the XPlorer Prize.
@@ Appendix A Related Work
Visual generative modeling has witnessed significant advancements in recent years. Recent explicit-likelihood approaches can be broadly categorized into diffusion-based models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Song et al., 2020b; Dhariwal & Nichol, 2021; Kingma et al., 2021; Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Karras et al., 2022; Bao et al., 2023; Peebles & Xie, 2023; Esser et al., 2024; Xie et al., 2024), auto-regressive models (Chen et al., 2020; Esser et al., 2021; Ramesh et al., 2021; Yu et al., 2021; Tian et al., 2024; Team, 2024; Sun et al., 2024; Ma et al., 2024; Zhang et al., 2024; Tang et al., 2024), and masked-prediction models (Chang et al., 2022; Yu et al., 2023a; Chang et al., 2023; Li et al., 2024; Yu et al., 2024; Fan et al., 2024). The introduction of guidance techniques has substantially improved the capabilities of these models. These include classifier guidance (Dhariwal & Nichol, 2021), classifier-free guidance (Ho & Salimans, 2022), energy guidance (Chung et al., 2022; Zhao et al., 2022; Lu et al., 2023; Song et al., 2023a), and various advanced guidance methods (Kim et al., 2022; Hong et al., 2023; Kynkäänniemi et al., 2024; Karras et al., 2024; Chung et al., 2024; Koulischer et al., ; Shenoy et al., 2024).
@@
To address the computational overhead introduced by classifier-free guidance (CFG), One widely used approach to remove CFG is guidance distillation (Meng et al., 2023), where a student model is trained to directly learn the output of a pre-trained teacher model that incorporates guidance. This idea of guidance distillation has been widely adopted in methods aimed at accelerating diffusion models (Luo et al., 2023; Yin et al., 2024; Lin et al., 2024; Zhou et al., 2024a). By integrating the teacher model’s guided outputs into the training process, these approaches achieve efficient few-step generation without guidance.
Alternative Methods for Building Guidance-free Models.
Recent studies in diffusion models show that perceptual losses (Lin & Yang, 2023), score-based distillation (Sauer et al., 2025a; Zhou et al., 2024c; Sauer et al., 2025b; Zhou et al., 2024b), and consistency models (Song et al., 2023b; Geng et al., 2024; Lu & Song, 2024) can also achieve comparable FID scores to CFG. For auto-regressive models, Condition Contrastive Alignment (Chen et al., 2024b) could enhance guidance-free performance through alignment (Rafailov et al., 2023; Chen et al., 2024a) in a self-contrastive manner.
@@ Appendix B Proof of Theorem 1
We first copy the training objective in Eq. (3.2) and Eq. (8).
@@
Given unlimited model capacity and training data, the optimal $\epsilon_{\theta^{*}}^{s}$ for optimizing Eq. (15) and Eq. (16) are the same. Both satisfy
@@
which is standard unconditional diffusion loss. According to Eq. (2) we have
@@
Then we prove for $\forall\beta\in(0,1]$, stopping the unconditional gradient does not change this optimal solution. Taking derivatives of ${\mathcal{L}}_{\theta}^{\text{practical}}$ we have:
@@
Since $2\mathbb{E}\beta]$ is a constant, this does not change the convergence point of ${\mathcal{L}}_{\theta}^{\text{raw}}$. The optimal unconditional solution for ${\mathcal{L}}_{\theta}^{\text{practical}}$ remains the same.
@@
Figure 11: Precision-Recall trade-off comparison for DiT and LlamaGen in fine-tuning experiments.
Figure 12: Precision-Recall trade-off comparison for DiT and LlamaGen in from-scratch-training experiments.
@@
We mainly apply GFT to fine-tune DiT-XL/2 (28 epochs, 2% of pretraining epochs) and train DiT-B/2 from scratch (80 epochs, following the original DiT paper’s settings (Peebles & Xie, 2023)). Since the DiT-B/2 pretraining checkpoint is not publicly available, we reproduce its pretraining experiment. For all experiments, we use a batch size of 256 and a learning rate of 1e41e-4. For DiT-XL/2 fine-tuning experiments, we employ a cosine-decay learning rate scheduler.
@@
We mainly apply GFT to fine-tune VAR-d30 models (15 epochs) or train VAR-d16 models from scratch (200 epochs). Batch size is 768. The initial learning rate is 1e51e-5 in fine-tuning experiments and 1e41e-4 in pretraining experiments. Following VAR (Tian et al., 2024), we employ a learning rate scheduler including a warmup and a linear decay process (minimal is 1% of the initial).
VAR by default adopts a pyramid CFG technique on predicted logits. The guidance scale 0 decreases linearly during the decoding process. Specifically, let n be the current decoding step index, and N be the total steps. The $n$-step guidance scale $s_{n}$ is
@@
We mainly apply GFT to fine-tune LlamaGen-3B models (15 epochs) or train LlamaGen-L models from scratch (300 epochs). For fine-tuning, the batch size is 256, and the learning rate is 2e42e-4. For pretraining, the batch size is 768, and the learning rate is 1e41e-4. We adopt a cosine-decay learning rate scheduler in all experiments.
@@
We apply GFT to MAR-B, including both fine-tuning (10 epochs) and training from scratch (800 epochs). We find the batch size crucial for MAR and use 2048 following the original paper. For fine-tuning, we employ a learning rate scheduler including a 5-epoch linear warmup to 8e48e-4 and a cosine decay process to 1e41e-4. For training from scratch, we employ a 100-epoch linear lr warmup to 8e48e-4, followed by a constant lr schedule, which is the same configuration as the original MAR pretraining.
@@
We apply GFT to fine-tune Stable Diffusion 1.5 (SD1.5) (Rombach et al., 2022) for 70,000 gradient updates with a batch size of 256 and constant learning rate of 1e51e-5 with 1,000 warmup steps. We disable conditioning dropout as we find it improves CLIP score.
For comparison, we also fine-tune SD1.5 using guidance distillation with a scale range from 1 to 14, while keeping other hyperparameters aligned with GFT.
For evaluation, following GigaGAN (Kang et al., 2023) and DMD (Yin et al., 2024), we generate images using 30K prompts from the COCO2014 (Lin et al., 2014) validation set, downsample them to 256×256, and compare with 40,504 real images from the same validation set. We use clean-FID (Parmar et al., 2022) to calculate FID and OpenCLIP-G (Ilharco et al., 2021; Cherti et al., 2023) to calculate CLIP score (Radford et al., 2021). All results are evaluated using 50 steps DPM-solver++ (Lu et al., 2022) with EMA models. The EMA decay rate is set to 0.9999.